A perpetuity pays $1 at the end of every year plus an additional $1 at the end of every second
year. The effective rate of interest is i = 5%. Find the present value of the perpetuity at time
0.
1
Expert's answer
2020-09-29T18:04:13-0400

solution


At the end of every year, the perpetuity pays $1. On every second year, an additional $1 is paid. The payment series formed is;



"(time,\ payment) =(1,\ 1), (2,\ 2),(3,\ 1),(4,\ 2),..."

Note that at the end of times "time = 1,3,5,7,..." , The perpetuity pays $1.

At the the end of times "time\ = 2,4,6,8,...." , The perpetuity pays $2.



"i= 0.05"



Assume (series 1) is the series of $1 payments and (series 2) be the series of$2 payments. Series 1 and 2 are level annuities. The present value can be obtained separately for each series and the solutions added together to get the present value of the perpetuity

series 1: (time = 1,3,5,7,...)


"payments,\ p=1" start at "time,\ t=1" and occur every "2\ years."


The present value:


"present\ value =p* \frac{1+i}{(1+i)^2-1}"


"=1* \frac{1.05}{(1.05)^2-1}=10.2439"


series 2: (time\ = 2,4,6,8,....)


"payments,\ p= 2" start at "time,\ t= 2" and occur every "2 \ years"


The present value:


"present\ value=p* \frac{1}{(1+i)^2-1}""=2* \frac{1}{(1.05)^2-1}=19.5122"

answer:

The present value of the perpetuity is the sum of the 2 series of payments

"=10.2439+19.5122""=29.7561"